OpenStudy (anonymous):

Why does lim x tends to 0 of sinx/tanx = 1 ? just can't see it

OpenStudy (anonymous):

Sin (x) / Tan (x) = Cos (x) since Tan(x) = Sin(x)/ Cos(x) and lim x-> 0 Cos(x) = 1 since Cos(0) = 1

OpenStudy (anonymous):

Is that a basic formula... Cosx = sinx/tanx??? Maybe I missed that Reading accidentally?

OpenStudy (anonymous):

Because tan(x) = sin(x)/cos(x), sin(x) divided by sin(x)/cos(x) is the same thing as sin(x)*cos(x)/sin(x). Cancel sin(x) and you're left with cos(x) with x approaching zero.

OpenStudy (anonymous):

Wonderful!!!Got it now! The answer always seems so simple after it's explained! Thank you again!